OpenStudy (mertsj):

f'(x) = (x^2-1)/x, f(1) = .5 and f(-1) = 0 Find f(x)

OpenStudy (anonymous):

\[f'(x)=\frac{x^2-1}{x}=x-\frac{1}{x}\] so \[f(x)=\frac{x^2}{2}-\ln(x)+C\] and there is no way that i can see that \[f(-1)\] exists. what am i missing?

OpenStudy (mertsj):

I thought the same thing but the textbook gives both of those values.

OpenStudy (anonymous):

\[f(x)=x ^{3}/3+lnx+c\]

OpenStudy (mertsj):

Actually, isn't the integral of 1/x ln|x|?